OpenStudy (anonymous):

can some one do this limit n->infinity (3^n -1)/n!

OpenStudy (anonymous):

so taking the ratio of two succesive terms: [(3*3^n - 1)/(n + 1)(n!)]*[n!/(3^n -1)] with a little algebraic manipulation you can see that this function is monotonically(sp?) decreasing. And by plugging in values, you can see that 0 < (3^n - 1)/(n!) < 2, so it must converge to a finite limit. Since it's always decreasing, it must approach zero. So lim n -> infty (3^n - 1)/(n!) = 0. ( i apologize for the messyness of this explanation, but i hope it's sufficient)

OpenStudy (zarkon):

not that it really matters...is it \[\frac{3^n-1}{n!}\] or \[\frac{3^{n-1}}{n!}\]

OpenStudy (anonymous):

it is the first

OpenStudy (zarkon):

well...look at \[\frac{3^n}{n!}\] \[=\frac{3\cdot 3\cdot 3\cdot 3\cdot 3\cdot 3\cdot 3\cdot 3\cdot 3\cdots 3}{1\cdot 2\cdot 3\cdot 4\cdot 5\cdot 6\cdot 7\cdot 8\cdot 9\cdots n}\le \frac{3}{n}\to 0\text{ as }n\to\infty\]

OpenStudy (anonymous):

@Zarkon This makes perfect sense, thank you very much.